OpenStudy (anonymous):

What would 7√x5 * 7√x5 be? please, I need help

OpenStudy (anonymous):

\[\sqrt[7]{x5} * \sqrt[7]{x5}\]

jhonyy9 (jhonyy9):

so do you can calcule without radicals x^5 *x^5 = ?

jhonyy9 (jhonyy9):

just because hope you know it sure that sqrt(x) *sqrt(x) = sqrt(x^2) yes ?

OpenStudy (anonymous):

oh, sorry, yeah, I got \[x ^{10/7}\]

jhonyy9 (jhonyy9):

yes you are right